A number is 9 more than another.The sum of the 2 number is 37.Find the two numbers.
Found 3 solutions by stanbon, Cromlix, Alan3354:
Answer by stanbon(75887)

A number is 9 more than another.The sum of the 2 number is 37.Find the two numbers.
---------------
x = y + 9
x + y = 37
-----
Substitute for "x" and solve for "y"::
y+9+y = 37
2y = 28
y = 14
---
x = y+9 = 14+9 = 23
-----------------------------
-------------
Cheers,
Stan H.

Answer by Cromlix(4381)

Hi there,
Make the original number = x
Other number = x + 9
x + (x + 9) = 37
x + x + 9 = 37
2x + 9 = 37
2x = 37 - 9
2x = 28
x = 14
Original number = 14
Other number = 23
Hope this helps :-)

Answer by Alan3354(69443)

x - y = 9
x + y = 37
------------ Add
2x = 46
x = 23
y = 14